RC meeting and Run Plan, Wednesday December 14th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Current RC: Eric King (turning over to Andrew)
Time: Daily at 3:30PM
Location: Counting House Conference Room

Remote Connection Details

Shift Schedule

Link to Schedule

Until Shutdown

All shifts are still full! :)

After Returning

There is a need for OWL shifters. Please encourage your students to take some owl shifts. A string of owl shifts are easier to do than doing one here and there.

  • Silviu has signed up for a few OWL shifts since I made the screensnip.


Accelerator Status

Very minor disruptions on the accelerator side. Two ~25 minute downs between SWING and OWL. Events: 17477, 17478

Current Status

Taking production data on 3-He

Shift Summaries

Moller Measurement

  • Moller measurment during SWING
20047 IN 10um Fe 86.11% 0.15%
20048 IN 11um Cu 0.07% 0.17%
20049 OUT 11um Cu -0.14% 0.16%
20050 OUT 10um Fe -86.46% 0.15%

Beam Current

We've been happily running at 40uA when taking production data.


Target polarization has been stable. List of the recent polarizations

Issues remaining

  • BB GEM APV unreachable and cannot be repaired at this time.

Short Term Run Plan

  • Continue production run on He3 target, 40 uA
  • Use GEnII-noSBSGEMs DAQ config
  • Use ps3=0, ps5=6, others =-1
  • BBCal threshold at -366 mV, HCal threshold at -19 mV
  • NMR measurement every 3 - 4 hours, close to 4 hours is better

Planned Accesses to the Hall


Special Requests


Notices in Effect for Shift Crew

  • Power down SBS, BB, and Detector HV before making an access to the Hall

Data Collection Status

  • 79 out of 141 days (~56.0%)
  • Total uncorrected accumulated charge on He3: 61.573 C out of 232C (26.54%)
    • 2-Pass: 13.467 C
    • 3-Pass: 48.106 C (Wednesday 8AM)